OpenStudy (anonymous):

Why k/0 is undefined (k is not equal to zero)

OpenStudy (anonymous):

I can explain this question.

OpenStudy (anonymous):

yes u may :)

OpenStudy (anonymous):

Let us suppose that k=2, than we get 2/0=x, so 2/x=0, isn't?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry, in my last reply, "than" should be"then". As 2/x=0, so we get 2=0*x. All number times 0 is 0, so 0*x=0. Then we get 2=0.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

2=0 dosen't mean undefined that's just a mathematical fallacy

OpenStudy (helder_edwin):

the argument shows that k/0 is undefined.

hartnn (hartnn):

actually 2/0 is infinity. but since infinity is not defined for you yet, you treat it as undefined. you know what is the notion of infinity ? i represents a very very large number. and k/0 has a very very small (infact =0) denominator, so overall the fraction will be very very large (infact infinity)

OpenStudy (anonymous):

There must be some logic behind it and i want the logic.

OpenStudy (anonymous):

yes i know infinity

hartnn (hartnn):

infinity represents... ****

OpenStudy (anonymous):

i is a complex number if i am not wrong

OpenStudy (anonymous):

Sorry I made a mistake. The denominator smaller, the value of fraction greater. So a mathematician had said k/0 should be infinity.

OpenStudy (helder_edwin):

k/0 "is" infinity only in the system of the extended real numbers.

OpenStudy (anonymous):

But why any pure logic behind it??

OpenStudy (anonymous):

My teacher couldn't explain it to me .

hartnn (hartnn):

thats the logic when denominator is small, fraction is large when denominator = 0, fraction = infinity

OpenStudy (anonymous):

ohk thanks all for answering

OpenStudy (anonymous):

But if the denominator is smaller than zero will it be greater than infinity???

OpenStudy (anonymous):

I know it can't be

hartnn (hartnn):

smaller than 0 >>>>>negative you mean values like -0.0000000001 ?? it would be near to -infinity then

hartnn (hartnn):

by >>>>>> , i meant 'means' and not greater than

OpenStudy (anonymous):

-1??? smaller than zero so should be greater than infinity?? but it is not

hartnn (hartnn):

what is smaller than 0 ?

hartnn (hartnn):

nothing is greater than infinity

OpenStudy (anonymous):

x<0 x{-1,-2,-3.................................................)

hartnn (hartnn):

yeah, so like k/-2 will not be greater than infinity right ?

OpenStudy (anonymous):